OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x^4 - 21x^2 - 100

OpenStudy (whpalmer4):

We have a polynomial with only real coefficients, so our complex roots must come in conjugate pairs. What is the conjugate of \(-2i\)?

OpenStudy (anonymous):

+2i?

OpenStudy (anonymous):

imaginary roots always come in pairs right? :)

OpenStudy (anonymous):

yea

OpenStudy (whpalmer4):

yes, you are correct. what is (x-2i)(x+2i)? We can divide that out of our polynomial to get a simpler polynomial to find the remaining roots.

OpenStudy (anonymous):

all i know so far is (x+2i)*(2-2i)*(x+a)*(x+b) =0 :P could be wrong tho.

OpenStudy (whpalmer4):

change that to (x+2i)(x-2i)(x+a)(x+b) = 0 and I'll agree with you

OpenStudy (jdoe0001):

\((x+2i)(x-2i) \implies x^2-(2i)^2 \implies x^2-4i^2 \implies x^2+4\)

OpenStudy (whpalmer4):

so, if we have \[x^2 + a^2 = 0\]\[x^2=-a^2\] \[x=\pm ai\] So that means we can skip multiplying and just go directly to \[(x-2i)(x+2i) = x^2+4\] Now let's divide our original polynomial by \(x^2+4\)

OpenStudy (whpalmer4):

a quick look suggests that it's going to be \(x^2-25\) because we can take out \(x^2(x^2+4) = x^4+4x^2\) and that leaves us with \(-25x^2-100\) which is just \(-25(x^2+4)\). What are the zeros of \(x^2-25\)?

OpenStudy (whpalmer4):

You should recognize that as a difference of squares...

OpenStudy (whpalmer4):

\[a^2-b^2 = (a-b)(a+b)\]

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

(x-25)(x+25)

OpenStudy (anonymous):

sorry!

OpenStudy (anonymous):

(x-5)(x+5)

OpenStudy (whpalmer4):

right. so the final two zeros are?

OpenStudy (anonymous):

+/- 5?

OpenStudy (whpalmer4):

yes. so the entire list of zeros of the polynomial is \[x = \{-2i,2i,-5,5\}\]